Use two equations in two variables to solve the application.
A merchant wants to mix peanuts worth $3 per pound with jelly beans worth $1.50 per pound to make 30 pounds of a mixture worth $2.10 per pound. How many pounds of each should he use? (find # of lbs used for peanuts and jelly beans).

Found 2 solutions by Alan3354, josgarithmetic:
Answer by Alan3354(69443)

A merchant wants to mix peanuts worth $3 per pound with jelly beans worth $1.50 per pound to make 30 pounds of a mixture worth $2.10 per pound. How many pounds of each should he use? (find # of lbs used for peanuts and jelly beans).
----------
p + j = 30 --- total mix
3p + 1.5j = 2.1*30 --- total cost
etc.

Answer by josgarithmetic(39620)